Mohave Inc. purchased land, building, and equipment from Laguna Corporation for a cash payment of $315,000. The estimated fair values of the assets are land $60,000, building $220,000, and equipment $80,000. At what amounts should each of the three assets be recorded?(Round final answer to 0 deciaml places, e.g. 5,275.)

                          Recorder Amount

Land

Building

Equipment
Answer #1

Total Asset Fair Value = Land + Building + Equipment

Total Asset Fair Value = 60000+220000+80000

Total Asset Fair Value = 360000

Recorder Amount

Land = 60000/360000 * 315000

Land = 52500

Building = 220000/360000 * 315000

Building = 192500

Equipment = 80000/360000 * 315000

Equipment = 70000

Answer

Recorder Amount
Land                              52,500
Building                            192,500
Equipment                              70,000